OpenStudy (anonymous):

Which of the following is another form of the equation 8^(2x-1)=32 A. 4^(2x-1)=8 B. 2^(6x-3)=2^5 C.4^(4x-1)=4^8 D. 1^(2x-1)=4

OpenStudy (whpalmer4):

\[8^{2x-1} = (2*2*2)^{2x-1}\]

OpenStudy (whpalmer4):

\[2^5=2*2*2*2*2\]

OpenStudy (anonymous):

all u do is reduce everything

OpenStudy (anonymous):

2^(6x-3)=2^5

OpenStudy (whpalmer4):

\[(ab)^n = a^nb^n\]

OpenStudy (anonymous):

Thank you everyone!